OpenStudy (cooltowl):

With 4.95 g of baking soda, how many moles of NaOH should be formed? Using stoichiometric calculations Reaction 1: 1 NaHCO3(s) → 1 NaOH(s) + 1 CO2(g)

OpenStudy (aaronq):

1. Convert the mass to moles 2. write a ratio of the chemical compounds of interest and their respective stoichiometric coefficients \(\sf \dfrac{moles~of~baking~soda}{baking~soda's~coefficient}=\dfrac{moles~of~NaOH}{NaOH's~coefficient}\) 3. solve for moles of NaOH use \(\sf moles=\dfrac{mass}{molar~mass}\) for 1. and 3.

OpenStudy (cooltowl):

would it look like this, .0589/1=.123759/1

OpenStudy (aaronq):

yes it would. for this example the ratio doesn't really matter because both coefficients are equal to 1, but this is how you solve all stoichiometry problems.

OpenStudy (aaronq):

wait, no, where did you get 0.123579?

OpenStudy (aaronq):

moles of NaOH should be your unknown variable in the ratio, what you are solving for.

OpenStudy (cooltowl):

.0589/1=?/1 .0589=?/1 How do I solve this because I keep on getting the wrong answer which is .0589

OpenStudy (aaronq):

it should be that, 0.058923 moles